OpenStudy (kendricklamar2014):

Find the first four terms of the sequence represented by the expression. 3n +5 A. 3, 8, 14, 20 B. 8, 11, 14, 17 C. 3, 6, 9, 12 D. 0, 8, 11, 14

OpenStudy (kendricklamar2014):

@paki

OpenStudy (anonymous):

i think it is C. 3, 6, 9, 12

hartnn (hartnn):

nopes, to find 1st term, plug in n=1 in 3n to find 2nd term, plug in n=2 in 3n +5 and so on

OpenStudy (anonymous):

sorry its not c and its b

OpenStudy (paki):

for n=1..... 3n +5 will be ==> 3(1)+5 = 3+5= 8 for n=2..... 3n +5 will be ==> 3(2)+5 = 6+5= 11 for n=3..... 3n +5 will be ==> 3(3)+5 = 9+5= 14 for n=4..... 3n +5 will be ==> 3(4)+5 = 12+5= 17 ..... so option B.... @KendrickLamar2014

OpenStudy (kendricklamar2014):

thanks

OpenStudy (paki):

pleasure